OpenStudy (anonymous):

solve system of equations: 3x-4y=25 x^2+y^2=25

OpenStudy (anonymous):

square equation (1) : 9x^2 - 16y^2 = 625 (Multiply equation (2) by 9 ) : 9x^2 + 9 y^2 = 225 solve : - 16y^2 - +9y^2 = 625 - 225 -25y^2 = 400 - y =\[\sqrt{400/25}\] - y = \[\sqrt{16}\] y = -4 if y = -4 then 3x = 25+4y = 25 + 4 ( -4) = 25-16 = 9 therefore x = 9 and y = -4 and it should be the right answer,,,, :)

OpenStudy (anonymous):

i didn't know you could square the first one that makes it simple then ! thanks

OpenStudy (anonymous):

you just need to cancel out either x or y... to obtain the answer.. i prefer to square the equation and multiply by 9 to the other equation rather than : \[x = (25 + 4y) \div 3\] and then plug in value of x into the equation : x^2 + y^2 = 25

OpenStudy (anonymous):

yea i did it like that and it was a mess .. squaring seems better in this case thank you!

OpenStudy (anonymous):

btw i made a mistake... when replacing y value into the equation: 3x = 25 + 4y 3x = 25 + 4 (-4) 3x = 9 x = 3 sorry..

OpenStudy (anonymous):

i get sqrt of -16 which will give me an imaginary how did you get a whole number

OpenStudy (anonymous):

i did it like : -y = sqrt of 16... just to avoid the imaginary so , y= -4

OpenStudy (anonymous):

is that 'legal' ?

OpenStudy (anonymous):

guess not :s i just did that to avoid - sqrt of 16 but let me try again another way ..

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

3x - 4y = 25 , x^2 + y^2 = 25 so : 3x -4y = x^2 + y^2 x^2 - 3x = - y^2 -4y x (x -3) = -y ( y + 4) ( x-3 ) = 0 or (y + 4 ) =0 x= 3 and y = -4

OpenStudy (anonymous):

that would mean that x is also 0 and y is also 0

OpenStudy (anonymous):

yep but you cant use this values because if x and y would be 0 it would mean that the curve does exist so it is rejected x = 0 y=0 ( Rejected )

OpenStudy (anonymous):

 this is the graph of x^2 + y^2 = 25